OpenStudy (anonymous):

How do you find out what Seventy percent of what number is 17.5?

jimthompson5910 (jim_thompson5910):

"Seventy percent of what number is 17.5" translates to 0.7x = 17.5 where x is the unknown number

jimthompson5910 (jim_thompson5910):

the 0.7 is the decimal form of 70%

OpenStudy (jdoe0001):

if we know that 17.5 is the 100% original number what is 70% of that? \(\begin{array}{ccllll} quantity&percentage \\\hline\\ 17.5&100\\ 70&x \end{array}\implies \cfrac{17.5}{70}=\cfrac{100}{x}\)

OpenStudy (jdoe0001):

hmmm I got my quantities swap off =) \(\begin{array}{ccllll} quantity&percentage \\\hline\\ 17.5&100\\ x&70 \end{array}\implies \cfrac{17.5}{x}=\cfrac{100}{70}j\)

OpenStudy (jdoe0001):

more typos \(\bf \begin{array}{ccllll} quantity&percentage \\\hline\\ 17.5&100\\ x&70 \end{array}\implies \cfrac{17.5}{x}=\cfrac{100}{70}\)

sammixboo (sammixboo):

First you have to change 70% into adecimal: 70=70/100=0.7 Then you basically divide 17.5 by 0.7: 17.5/0.7=?

OpenStudy (anonymous):

@sammixboo thanks :-) it helped a lot!

sammixboo (sammixboo):

No problem.. What answer did you get?

OpenStudy (anonymous):

25 :) @sammixboo

sammixboo (sammixboo):

That's correct! Glad I could help with your problem :D \(\it\large\color{#00FF17}-\color{#00FF5C}S\color{#00FF8A}a\color{#00FFB8}m\color{#00FFE6}m\color{#00E6FF}i\color{#00CFFF}♥\)

sammixboo (sammixboo):

If you need anymore help just tag my name in the comments of your queston and I will be there asap! \(\it\large\color{#00FF17}-\color{#00FF5C}S\color{#00FF8A}a\color{#00FFB8}m\color{#00FFE6}m\color{#00E6FF}i\color{#00CFFF}♥\)

OpenStudy (anonymous):

@sammixboo I will, thanks for the help I really appreciate it :D

sammixboo (sammixboo):

No problem ;) \(\it\large\color{#00FF17}-\color{#00FF5C}S\color{#00FF8A}a\color{#00FFB8}m\color{#00FFE6}m\color{#00E6FF}i\color{#00CFFF}♥\)